From a sack of fruit containing ​3 bananas, 3 oranges​, and 3 apples ​, a random sample of 4 pieces of fruit is selected. Suppose X is the number of bananas and Y is the number of apples in the sample.



​(a) Find the joint probability distribution of X and Y.



​(b) Find ​P[(X,Y)​A], where A is the region that is given by (x,y)



Question content area bottom



Part 1



​(a) Complete the joint probability distribution below.



​(Type integers or simplified​ fractions.)



1
Expert's answer
2021-12-23T06:22:14-0500

Solution:

(a):

"P(X=x,Y=y)=\dfrac{^3C_x\ ^3C_y\ ^3C_{4-x-y}}{^9C_4}; x=0,1,2,3,y=0,1,2,3"

So, "P(X=0,y=1)=\dfrac{^3C_0\ ^3C_1\ ^3C_{4-0-1}}{^9C_4}=0.0238"

"P(X=1,y=1)=\dfrac{^3C_1\ ^3C_1\ ^3C_{4-1-1}}{^9C_4}=0.2143"

and so on till "P(X=4,y=4)=\dfrac{^3C_4\ ^3C_4\ ^3C_{4-4-4}}{^9C_4}=0"


(b):

"P[(x,y),x+y\le1]=f(0,0)+f(0,1)+f(1,0)
\\=0+0.0238+0.0238
\\=0.0476"